OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). If you could walk me through to the answer with steps that'd be great!!! THANK YOU

OpenStudy (anonymous):

A. y = -one thirdx - 6 B. y = -one thirdx - 14 C. y = 3x - 6 D. y = 3x - 14

OpenStudy (texaschic101):

a parallel line will have the same slope. So we have to find the slope in the equation 3x - y = -3. To do this, we will put the equation in y = mx + b form, where the number in the m position is the slope. 3x - y = -3 -- subtract 3x from both sides -y = -3x - 3 -- now multiply by -1 to make y positive y = 3x + 3 the number in the m position is 3, therefore the slope is 3. You still with me ?

OpenStudy (texaschic101):

we will be using the same slope, 3. Now we will use y = mx + b slope(m) = 3 (4,-2) x = 4 and y = -2 we will now sub and solve for b y = mx + b -2 = 3(4) + b -2 = 12 + b -- subtract 12 from both sides -2 - 12 = b -14 = b so our parallel equation is : y = 3x - 14 do you have any questions ?

OpenStudy (anonymous):

Yes how did you get -14?

OpenStudy (texaschic101):

I used my points (4,-2) and my slope 3 and subbed it into y = mx + b form to solve for b, the y intercept. If you will look at my post that I posted, it shows how I got -14. Do you still not understand ? Because there is another way to do this, if you want me to show you

OpenStudy (texaschic101):

do you need me to show you the other way ?

OpenStudy (anonymous):

If it isn't too much trouble, yes.

OpenStudy (texaschic101):

you will still need to find the slope of the equation, which is 3. We now have the slope and a set of points (4,-2). we can use the point-slope formula : y - y1 = m(x - x1) slope(m) = 3 (4,-2) x1 = 4 and y1 = -2 now we will sub this into the formula y - y1 = m(x - x1) y - (-2) = 3(x - 4) -- distribute through the parenthesis y + 2 = 3x - 12 -- subtract 2 from both sides y = 3x - 12 - 2 -- combine like terms y = 3x - 14 either way you do it, you will arrive at the same answer

OpenStudy (texaschic101):

if you still have questions, just ask and I will explain the best I can :)